Thirty pounds of dried fruit mix contained apples worth $3.15 a pound and pears worth $4.25 a pound. How many pounds of each did the mix contain if the value of the mix was $107.70?
Answer by josmiceli(19441)

x = pound of apples
30 - x = pounds of pears
107.70 = 3.15x + (30 - x)4.25
expand
107.70 = 3.15x + 127.50 - 4.25x
combine terms
-19.80 = - 1.10x
x = 18
30 - x = 12
18 pounds of apples and 12 pounds of pears
check
107.70 = 3.15(18) + (12)4.25
107.70 = 56.70 +51.00
107.70 = 107.70